• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jmammei

    Log entry time 15:50:49 on June17,2010

    Entry number 327082

    This entry is a followup to: 327060

    keyword=N2 ice flashed off!

    Ed went down to turn on the pumps so that we could try to get to 1e-4 before 6pm so that we could still do the 70uA test. I went upstairs to talk to Paul. We were considering abandoning the warm-up because it wasn't working - when I left the counting house the target ladder was still covered in fuzz. Ed looked at the target right before coming upstairs and still saw it. He turned on the normal vacuum gauge, and the pressure was dropping, as shown in the attached figure.

    It must have been N2 ice because by the time we both returned to the counting house it was gone. So we will continue to try to pump down in time to do the 70uA test at 6pm. If all goes well we may be able to run with high current overnight. Keep your fingers crossed!


    FIGURE 1